OpenStudy (anonymous):

Year 3, F(3) = 150(1.02)^3 = 159. Year 5, F(5) = 150(1.02)^5 =166 Year 7, F(7) = 150(1.02)^7 = 172 What would be the average rate of change between year 3 and 5 and the rate of change between 5 and 7.

OpenStudy (anonymous):

So for years 3 and 5 this is what i have so far.\[\frac{ f(5) - f(3) }{b - a }\]

OpenStudy (anonymous):

is this correct or not?

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

Thats not correct but could someone please help meee

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

from 3 to 5, you replace b with 5 and a with 3 \[\Large \frac{ f(5) - f(3) }{b - a } = \frac{ f(5) - f(3) }{5-3 } = ??\]

OpenStudy (anonymous):

So thats what you would do for the top half of the formula what would you do for the bottom half?

OpenStudy (anonymous):

5-3 is 2 so if both of them were 5-3 it would end up being 2/2 which would be 1

jimthompson5910 (jim_thompson5910):

just do 5-3 = 2 for the bottom half

jimthompson5910 (jim_thompson5910):

f(5)-f(3) is not equal to 2 though

jimthompson5910 (jim_thompson5910):

it's equal to the difference in the outputs you got at the top

OpenStudy (anonymous):

What is it equal to?

jimthompson5910 (jim_thompson5910):

Year 3, F(3) = 150(1.02)^3 = 159. Year 5, F(5) = 150(1.02)^5 =166

jimthompson5910 (jim_thompson5910):

assuming you did that correctly

OpenStudy (anonymous):

Ohhh i see

OpenStudy (anonymous):

7/2

OpenStudy (anonymous):

So now that i have 7/2 what would i do next

OpenStudy (anonymous):

Or would that be it? is the question complete?

jimthompson5910 (jim_thompson5910):

yeah that is the approximate avg rate of change

OpenStudy (anonymous):

Okay so would i just do the same for year 5 through 7?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

Okay if i have any other questions ill ask you!!